OpenStudy (anonymous):

How do you find the cross products to determine whether the following proportions are true or false ? 1.) 2/3=20/27 2.) 110/18=16/3

Directrix (directrix):

You use the Means-Extremes property. a/b = c/d IF a*d = b*c

Directrix (directrix):

1.) 2/3=20/27 Does 2*27 = 3*20? If so, the proportion is correct. If not, the proportion is false.

OpenStudy (anonymous):

Oh thank you !

Directrix (directrix):

Means-Extremes Property and Cross Products Property are the same.